Kevin is 15 years older than Jessica. Kevin and Jessica first met two years ago. Eighteen years ago, Kevin was 4 times as old as Jessica.

Answer by TimothyLamb(4379)

x = kevin age now
y = jessica age now
---
x = y + 15
x - 18 = 4(y - 18)
---
x = y + 15
x - 18 = 4y - 72
---
x = y + 15
x - 4y = -54
---
put the system of linear equations into standard form
---
x - y = 15
x - 4y = -54
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = kevin age now = 38
y = jessica age now = 23
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php